# Generative AI for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

### v1

@@ 6.2.5 Interactive Wildfire Management through Explainable AI and Cognitive Digital Twins
Energy and Environmental Generative AI model development has a significant environmental impact due to high energy consumption and carbon emissions from large-scale training. Mitigation strategies include model distillation, sparsity optimization, transfer learning, and adopting energy-efficient ”green AI” practices.
@@ 6.3.1 Computational Challenges
Training large-scale generative AI models—particularly diffusion models, transformers, and other autoregressive architectures—presents substantial computational challenges. These models typically require high-resolution spatial-temporal data spanning vast geographic regions and extended fire seasons, leading to massive, memory-intensive datasets. As emphasized by Manduchi et al. (Manduchi et al., 2024), training such models at scale demands access to powerful GPUs or TPUs, careful memory management, and optimization strategies to maintain training stability across distributed systems.
Moreover, inference with diffusion models often involves iterative denoising steps, while transformer-based language models follow a sequential token generation process, both of which result in high latency. This is particularly problematic in real-time or near-real-time wildfire forecasting applications where rapid decision-making is essential. Even recent improvements like FlashAttention and latent-space diffusion training (Manduchi et al., 2024; Bandi et al., 2023) only partially mitigate these bottlenecks. Additionally, robust training requires extensive data augmentation, domain adaptation, and fine-tuning, which further increases the computational burden (Bandi et al., 2023).
To improve computational efficiency, emerging strategies such as model quantization, low-bit training, distillation, and lossy latent-space modeling have shown promise (Manduchi et al., 2024). However, these approaches often come with trade-offs in terms of accuracy, robustness, or interpretability—making them challenging to adopt in high-stakes domains such as environmental hazard forecasting and disaster response.
@@ 6.3.2 Evaluation Challenges
Evaluating the outputs of generative AI models in wildfire prediction poses significant methodological challenges, particularly due to the lack of standardized, interpretable, and domain-specific evaluation frameworks. Unlike classification or regression tasks, where performance can be quantified using well-established metrics, the assessment of synthetic wildfire scenarios—especially those generated under creative or hypothetical conditions—remains highly subjective and context-dependent (Bandi et al., 2023; Manduchi et al., 2024). Widely used metrics such as Fréchet Inception Distance (FID) or Inception Score (IS), originally developed for image synthesis, are ill-suited for evaluating spatiotemporal fidelity, physical realism, or consistency with known fire behavior dynamics in environmental simulations (Bandi et al., 2023; Manduchi et al., 2024).
Recent studies emphasize the urgent need for developing “domain-specific evaluation metrics” tailored to complex generative tasks. These include assessing spatial spread accuracy, temporal progression, and physical plausibility relative to empirical wildfire data and simulation-based fire behavior models (Fui-Hoon Nah et al., 2023). Additionally, generative AI models frequently suffer from pathologies such as “mode collapse, sample hallucination, and memorization”, which compromise their generalizability and reliability when trained on sparse or biased fire datasets (Manduchi et al., 2024). These issues are particularly critical in “high-stakes applications” such as real-time emergency response or predictive decision support, where misleading outputs can result in dangerous operational consequences.
The evaluation challenge is further exacerbated by the absence of ground truth for rare or extreme wildfire events, which limits model validation using traditional benchmarks. As Manduchi et al. (2024) and Sun et al. (2024) argue, current benchmarks fail to capture domain-specific constraints and interpretability, especially in dynamic, real-world environments. To address these shortcomings, researchers have advocated for the integration of “human-in-the-loop evaluation”, “uncertainty quantification”, and “physics-informed priors” into both model training and assessment pipelines (Yan et al., 2024; Fui-Hoon Nah et al., 2023). Furthermore, a growing body of literature calls for “hybrid metrics” that combine statistical quality measures with expert-driven assessments and simulation-based validations, offering a more holistic evaluation of generative outputs (Bandi et al., 2023; Manduchi et al., 2024; Fui-Hoon Nah et al., 2023).
@@ 6.3.3 Energy and Environmental Challenges
One significant but often under-discussed challenge in developing generative AI models for wildfire prediction lies in their substantial energy consumption and environmental footprint. Training foundation models—particularly those based on large-scale transformer architectures—requires extensive computational resources, often involving thousands of GPU hours and high-performance computing clusters (Zhou et al., 2024). This process results in considerable electricity usage, much of which is still powered by fossil fuels in many regions, leading to high levels of carbon emissions. As highlighted by Myers et al. (2024), the environmental cost of training a single large language model can surpass the annual carbon footprint of several individuals. While these models offer valuable capabilities for simulating fire dynamics and generating proactive mitigation strategies, their development raises important concerns regarding sustainability and ecological responsibility, especially in the context of wildfire prediction, where climate change and ecosystem degradation are already pressing issues.
To mitigate these impacts, it is essential to explore strategies such as model distillation, sparsity optimization, transfer learning from pre-trained models, and leveraging green AI principles that emphasize energy-efficient training and inference (Salehi and Schmeink, 2023; Barbierato and Gatti, 2024). Integrating these strategies not only aligns the development of AI models with climate-conscious values but also ensures that the tools designed to protect the environment do not inadvertently contribute to its degradation.
@@ 7 Conclusion
This paper advocates for a paradigm shift in wildfire prediction—from traditional physics-based and deep learning models to generative AI-driven approaches capable of real-time, high-resolution, and multimodal fire simulation. Through a comprehensive literature review and LLM-assisted bibliometric analysis, we highlight the strengths of emerging generative models—including VAEs, GANs, Transformers, Diffusion Models, and Flow-Based Models—in learning complex wildfire dynamics from heterogeneous, sparse, and incomplete datasets. These models offer enhanced scalability, probabilistic forecasting, and the capacity to simulate fire spread across both 2D and 3D spatial domains, effectively capturing large-scale patterns embedded in observational and simulated data. We further explore three key innovations: (1) the development of interdisciplinary AI foundation models that integrate urban and environmental data; (2) conversational multi-agent systems powered by LLMs and Retrieval-Augmented Generation (RAG) for user-centered fire intelligence; and (3) lightweight generative models optimized for edge devices, enabling on-site predictions in real time, even under limited connectivity. While promising, these approaches face critical challenges in multimodal data fusion, model interpretability, energy efficiency, and edge deployment. Overcoming these barriers is essential for realizing scalable, trustworthy, and sustainable AI systems for next-generation wildfire management and response.

### v2

@@ 6.1.5 Trustworthy Wildfire Prediction via Explainable AI and Interactive Visual Analytics
Stochastic Nature of Wildfire Behavior Wildfire spread exhibits inherent unpredictability due to sensitivity to initial conditions, microclimate fluctuations, turbulent convection, and ember-driven spotting, which no model can fully eliminate. Employ ensemble simulations, scenario-based planning, and robust uncertainty quantification to capture probabilistic ranges of outcomes, supporting risk-informed decision-making rather than exact forecasts.
@@
Energy and Environmental Generative AI model development has a significant environmental impact due to high energy consumption and carbon emissions from large-scale training. Mitigation strategies include model distillation, sparsity optimization, transfer learning, and adopting energy-efficient ”green AI” practices Salehi and Schmeink (2023); Barbierato and Gatti (2024).
@@ 6.2 Challenges and Potential Solutions
6.2.1 Stochasticity Challenges in Fire Prediction
While advances in computational modeling and data-driven learning have markedly improved our ability to simulate wildfire spread, it is critical to acknowledge the inherent stochasticity and irreducible unpredictability that characterize wildfire behavior Sullivan (2009); Morvan (2011). Unlike controlled physical systems, wildfires evolve within highly dynamic and often chaotic natural environments, where small variations in initial conditions—such as localized wind gusts, abrupt changes in humidity, or micro-scale fuel heterogeneity—can lead to dramatically different outcomes over time Tedim et al. (2018); Hilton et al. (2015). This sensitivity to initial and boundary conditions underscores a fundamental limitation: even the most sophisticated models, whether physics-based or generative, cannot entirely eliminate uncertainty in predicting fire evolution.
Moreover, wildfires are influenced by a multitude of coupled processes across scales, including turbulent convection, ember lofting and spotting, and rapid transitions in combustion regimes, which are intrinsically probabilistic and only partially observable Sullivan (2009). For example, fine-scale wind vortices can transport embers far beyond modeled fire perimeters, triggering new ignitions in a manner that defies deterministic forecasting. Similarly, fuel moisture content can fluctuate over short distances due to microclimatic effects, introducing additional variability that is difficult to capture with static or coarse-grained input data Hilton et al. (2015).
Consequently, while generative AI offers powerful new avenues for representing complex spread dynamics and for sampling from learned distributions that mirror observed variability, it does not circumvent the fundamental unpredictability embedded in wildfire phenomena. Instead, these models should be viewed as tools to better characterize the probabilistic ranges of likely outcomes and to improve computational efficiency, rather than to produce more precise point forecasts. This perspective emphasizes the need for ensemble simulations, scenario-based planning, and robust uncertainty quantification as integral components of any wildfire prediction framework Rochoux et al. (2018); Hilton et al. (2015). By explicitly accounting for the stochastic nature of fire behavior, researchers and practitioners can make more informed decisions that appropriately balance forecast precision with the intrinsic volatility of wildfire systems.
@@ 6.2.2 Computational Challenges
Training large-scale generative AI models—particularly diffusion models, transformers, and other autoregressive architectures—presents substantial computational challenges. These models typically require high-resolution spatial-temporal data spanning vast geographic regions and extended fire seasons, leading to massive, memory-intensive datasets. As emphasized by Manduchi et al. Manduchi et al. (2024), training such models at scale demands access to powerful GPUs or TPUs, careful memory management, and optimization strategies to maintain training stability across distributed systems.
Moreover, inference with diffusion models often involves iterative denoising steps, while transformer-based language models follow a sequential token generation process, both of which result in high latency. This is particularly problematic in real-time or near-real-time wildfire forecasting applications where rapid decision-making is essential. Even recent improvements like FlashAttention and latent-space diffusion training Manduchi et al. (2024); Bandi et al. (2023) only partially mitigate these bottlenecks. Additionally, robust training requires extensive data augmentation, domain adaptation, and fine-tuning, which further increases the computational burden Bandi et al. (2023).
To improve computational efficiency, emerging strategies such as model quantization, low-bit training, distillation, and lossy latent-space modeling have shown promise Manduchi et al. (2024). However, these approaches often come with trade-offs in terms of accuracy, robustness, or interpretability—making them challenging to adopt in high-stakes domains such as environmental hazard forecasting and disaster response.
@@ 6.2.3 Evaluation Challenges
Evaluating the outputs of generative AI models in wildfire prediction poses significant methodological challenges, particularly due to the lack of standardized, interpretable, and domain-specific evaluation frameworks. Unlike classification or regression tasks, where performance can be quantified using well-established metrics, the assessment of synthetic wildfire scenarios—especially those generated under creative or hypothetical conditions—remains highly subjective and context-dependent Bandi et al. (2023); Manduchi et al. (2024). Widely used metrics such as Fréchet Inception Distance (FID) or Inception Score (IS), originally developed for image synthesis, are ill-suited for evaluating spatiotemporal fidelity, physical realism, or consistency with known fire behavior dynamics in environmental simulations Bandi et al. (2023); Manduchi et al. (2024).
Recent studies emphasize the urgent need for developing “domain-specific evaluation metrics” tailored to complex generative tasks. These include assessing spatial spread accuracy, temporal progression, and physical plausibility relative to empirical wildfire data and simulation-based fire behavior models Fui-Hoon Nah et al. (2023). Additionally, generative AI models frequently suffer from pathologies such as “mode collapse, sample hallucination, and memorization”, which compromise their generalizability and reliability when trained on sparse or biased fire datasets Manduchi et al. (2024). These issues are particularly critical in “high-stakes applications” such as real-time emergency response or predictive decision support, where misleading outputs can result in dangerous operational consequences.
The evaluation challenge is further exacerbated by the absence of ground truth for rare or extreme wildfire events, which limits model validation using traditional benchmarks. As Manduchi et al. (2024) and Sun et al. (2024) argue, current benchmarks fail to capture domain-specific constraints and interpretability, especially in dynamic, real-world environments. To address these shortcomings, researchers have advocated for the integration of “human-in-the-loop evaluation”, “uncertainty quantification”, and “physics-informed priors” into both model training and assessment pipelines Yan et al. (2024); Fui-Hoon Nah et al. (2023). Furthermore, a growing body of literature calls for “hybrid metrics” that combine statistical quality measures with expert-driven assessments and simulation-based validations, offering a more holistic evaluation of generative outputs Bandi et al. (2023); Manduchi et al. (2024); Fui-Hoon Nah et al. (2023).
@@ 6.2.4 Energy and Environmental Challenges
One significant but often under-discussed challenge in developing generative AI models for wildfire prediction lies in their substantial energy consumption and environmental footprint. Training foundation models—particularly those based on large-scale transformer architectures—requires extensive computational resources, often involving thousands of GPU hours and high-performance computing clusters Zhou et al. (2024). This process results in considerable electricity usage, much of which is still powered by fossil fuels in many regions, leading to high levels of carbon emissions. As highlighted by Myers et al. (2024), the environmental cost of training a single large language model can surpass the annual carbon footprint of several individuals. While these models offer valuable capabilities for simulating fire dynamics and generating proactive mitigation strategies, their development raises important concerns regarding sustainability and ecological responsibility, especially in the context of wildfire prediction, where climate change and ecosystem degradation are already pressing issues.
To mitigate these impacts, it is essential to explore strategies such as model distillation, sparsity optimization, transfer learning from pre-trained models, and leveraging green AI principles that emphasize energy-efficient training and inference Salehi and Schmeink (2023); Barbierato and Gatti (2024). Integrating these strategies not only aligns the development of AI models with climate-conscious values but also ensures that the tools designed to protect the environment do not inadvertently contribute to its degradation.
@@ 7 Conclusion
This review systematically examined the emerging role of generative AI models—including VAEs, GANs, Transformers, and diffusion architectures—in advancing bushfire prediction, monitoring, and risk mapping. While only 11 studies were identified that specifically explore these approaches for wildfire science, their collective findings demonstrate that generative AI can achieve prediction accuracies near 90% and offer computational efficiencies far surpassing traditional methods. Despite these advances, the use of generative AI in this domain remains nascent compared to its widespread adoption in other scientific and engineering fields.
Building on these observations, we proposed several promising research directions to shape the next generation of wildfire decision support systems. These include developing unified multimodal frameworks that seamlessly integrate 2D and 3D data, designing conversational agentic AI systems to deliver interactive, real-time wildfire intelligence, training interdisciplinary AI foundation models, enabling edge-based scenario generation on mobile and IoT devices, and advancing explainable AI interfaces for improved transparency and trust.
Addressing these frontiers will require tackling notable challenges such as managing inherent wildfire stochasticity, reducing computational and energy footprints, and establishing rigorous, domain-specific evaluation standards. By harnessing the unique strengths of generative AI while embedding principles of explainability, sustainability, and interdisciplinary integration, future research can build robust, adaptive, and human-centered systems that significantly enhance wildfire prediction and emergency management capabilities.